OpenStudy (anonymous):

From which graph can you best find the solution to the equation (2/x)+(15/x^2)=1?

OpenStudy (anonymous):

Here are the options A, B, C, and D I keep getting none of these, I don't understand? The solution to the equation is -3,5

OpenStudy (anonymous):

I keep getting this graph like the one on this website, but it says find best solution which one is the best? they all don't seem right?? http://www.wolframalpha.com/input/?i=solution+2%2Fx%2B15%2Fx%5E2%3D1

OpenStudy (solomonzelman):

Here, so your equation is \(\huge\color{blue}{ \frac{2}{x} + \frac{15}{x^2} =1}\)

OpenStudy (anonymous):

Yes!

OpenStudy (solomonzelman):

Ok, I'll type it out step by step. Focus.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{2}{x}+\frac{15}{x^2} =\frac{1}{1} } \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{2x}{x^2}+\frac{15}{x^2} =\frac{1}{1} } \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{2x+15}{x^2} =\frac{1}{1} } \]

OpenStudy (solomonzelman):

Cross multiply. \[\huge\color{blue}{ 2x+15 =x^2} \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ -x^2+2x+15 =0} \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ x^2-2x-15 =0} \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ x^2-2x =15} \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ x^2-2x+1 =15+1} \] \[\huge\color{blue}{ x^2-2x+1 =16} \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ (x-1)(x-1)=16} \]\[\huge\color{blue}{ x-1=± 4} \]\[\huge\color{blue}{ x=1± 4} ~~~~~~~~~x=~5~~~~or~~-3\]

OpenStudy (anonymous):

Right :) same solution i got! but how do u relate this to my graph dilemma?

OpenStudy (solomonzelman):

I can't really say, because the attachment doesn't work for me.

OpenStudy (anonymous):

Which on the doc. doesn't work?? graphABCD i can draw the four options!

OpenStudy (solomonzelman):

Yeah, the graph ABCD doesn't work.

OpenStudy (anonymous):

 Option A